The ages of Grandad, Uncle John and Tom add up to 105 years. Grandad is twice as old as Uncle John, and Uncle John is twice as old as Tom. Grandad is how old? Uncle John is how old? Tom is how old?
What is the best way to approach these type of questions?

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
"The ages of Grandad, Uncle John and Tom add up to 105 years."
g + j + t = 105
"Grandad is twice as old as Uncle John,"
g = 2j
"and Uncle John is twice as old as Tom."
j = 2t
divide both sides by
t = .5j
:
Replace g and t in the 1st equation
2j + j + .5j = 105
3.5j = 105
j = 105/3.5
j = 30 yr is John;s age
:
You should be able to find the other ages, ensure that they add up to 105
Grandad is how old? Uncle John is how old? Tom is how old?